### v1

# Non-Equilibrium Aspects of Fission Dynamics within the Time Dependent Density Functional Theory

Abstract
In this report we will cover in some details only our preliminary results for the induced fission of odd-mass and odd-odd nuclei within a time-dependent density functional (TDDFT) extended to superfluid systems. In the talk we have covered three topics; i) Scission neutrons, ii) Induced fission of odd-mass and odd-odd nuclei, and iii) Nontrivial aspects of fission dynamics, which we discuss here briefly.
1 Introduction
The scission neutrons have been conjectured to exist by Bohr and Wheeler Bohr:1939 and more than 50 years after discovery of nuclear fission Wagemans Wagemans:1991 stated (see his introductory remarks) that they most likely do not exist. Wagemans’s remarks are notable in another respect also, he never refers to Meitner and Frisch breakthrough paper Meitner:1939 and instead gives all the credit to Bohr and Wheeler Bohr:1939 for explaining the nature of nuclear fission, even though Bohr and Wheeler refer to Meitner and Frisch in their paper and even the name nuclear fission was coined by Meitner and Frisch. The experimental search and various theoretical models predicted the existence of scission neutrons with contradicting conclusions and properties in the following years, as either dominating the number of emitted neutrons by fission fragments (FFs) or being almost absent. In a very recent paper Abdurrahman:2024 we have reported the results of our most comprehensive theoretical analysis of scission neutrons within the Time-Dependent Density Functional Theory (TDDFT) extended to superfluid fermionic systems and predicted an extensive range of their properties. We invite the reader to consult this publication Abdurrahman:2024 , including the quite extensive supplement.
In another relatively recent publication Bulgac:2024a we have presented several new aspects of the fission process, which were never discussed in literature and this topic will not be presented here in more detail. In particular we have presented evidence that the fission dynamics from saddle-to-scission has an intrinsically non-Markovian character and consequently various previous attempts to describe this process using various stochastic or kinetic approaches are inadequate. Moreover, since in nuclear large amplitude collective dynamics (LACM) the occupation probabilities of single-particle states with large momenta have very long tails $\propto C/p^{4}$, where $p$ is the local single-particle momentum, a significant repopulation of the quasiparticle states occurs during the fission dynamics and theoretical modeling with a reduced basis set of single-particle states, such as (TD)BCS+TDHF or even TDDFT with a limited number of quasiparticle states lead to inaccurate predictions of the fission fragments (FFs), or even fail to lead to fission.
The next section will cover only a broad overview of our new results concerning the induced fission of odd-mass and odd-odd nuclei, which have not been submitted yet for publication and are still under analysis.
Figure 1: In the upper row left panel we display the time difference to full spatial separation of the two FFs between the case of an even-even and odd-mass nuclei. The odd-mass system is treated with an exact odd-neutron and even-proton numbers respectively. The odd-mass nucleus is treated as having an even-neutron and an even-proton numbers, with the average $N=147$ and $Z=92$ respectively for 239U and then a quasiparticle state is excited as described in Refs. Ring:2004 ; Bertsch:2009 . In the upper row right panel we show several TDDFT trajectories for the fission of an odd-nucleon nucleus with solid lines and for reference a TDDFT trajectory for an even-even nucleus with average $N=147$ and $Z=92$, with a dashed-thin solid line. Unlike the TDDFT trajectories, in which typically the quadrupole moment $Q_{20}$ typically increases in time in the case of fission of odd-nucleon systems the nucleus encounters a “bumpier” potential energy surface in the $(Q_{20},Q_{30})$ plane towards scission and the system it gets “confused” following a much more "challenging obstacle course," see upper row right panel. In the second row we show the occupation probability of the initial odd-nucleon state as a function of time. For time $t=0$ the occupation probability was obtained by constructing the corresponding canonical occupation probabilities Bulgac:2023 . For $t>0$ apart from one canonical quasiparticle state with occupation probability 1, the rest of the quasiparticle states are all double occupied with probabilities less than 1.
2 Induced fission of odd-mass and odd-odd nuclei
Ever since the development of a microscopic description of pairing correlations by Bardeen et al. Bardeen:1957 and the recognition that they are also most likely crucial in explaining nuclear properties Bohr:1958 , the treatments of fermion systems with odd numbers of particles have relied on the so called Pauli Blocking approximation. The extra odd nucleon carries a non-vanishing angular momentum and as a result the nuclear mean field brakes time-reversal invariance and the entire nucleus as well has a negative time-parity. From the theoretical point of view this is a rather inconvenient feature, which brings in quite a number of technical difficulties in the treatment of nuclear fission, and which forced most theoretical treatments to adopt a number of approximations and assumptions. In nuclear LACM the Pauli blocking approximation has been used almost always, see the most recent and detailed studies of the fission of odd-mass nuclei, where Pauli blocking in conjunction along with the equal filling approximation Perez-Martin:2008 have been used Schunck:2023 ; Pore:2024 .
We will present briefly here an overview of our results in the study of the fission of odd nuclei 239U (23 trajectories), 241Pu (42 trajectories), 243Pu (13 trajectories), and the odd-odd nucleus 238Np (35 trajectories), where in each case the number of trajectories describe various initial nuclear shape deformations characterized by quadruple and octupole momenta $(Q_{0},Q_{30})$ and various odd quasiparticle states of the respective fissioning nucleus, constructed using the prescription outlined in Refs. Ring:2004 ; Bertsch:2009 ; Schunck:2023 ; Pore:2024 . In total we have accumulated 113 trajectories for these nuclei, all obtained within the TDDFT extended to superfluid systems, a number comparable (or maybe even bigger) with perhaps of all of the rest of the world TDDFT fission trajectories obtained for even-even nuclei by other authors since 2015. The simulations have been performed in a simulation box $30^{2}\times 60$ and following the dynamics for up to about 60,000 times steps or longer in some cases. These simulations were performed with the energy density functional SeaLL1 Bulgac:2018 by evolving $0002\times 2\times 4\times 30^{2}\times 60=864,$ 000 complex, coupled, nonlinear PDEs in 3D+time, see Refs Bulgac:2019c ; Bulgac:2020 ; Bulgac:2024a for details.
Overall the FF properties in the induced fission of even-even, odd-mass, and odd-odd nuclei are very similar. Often there is a difference in the times the two FFs reach scission, when the corresponding initial quasiparticle state has a relatively low excitation energy, typically smaller than 1 MeV, see Fig. 1. In that case the TDDFT trajectory is significantly more complicated as the system spends quite a lot of time ’deciding’ which path to follow on the potential energy surface. Such a situation has been encountered for some almost symmetric TDDFT trajectories Bulgac:2019c ; Bulgac:2020 in the case of fission even-even compound nuclei.
The most unexpected feature is however the violation of the Pauli blocking approximation, see Fig. 1, used for more than 6 decades in nuclear physics. During the time evolution, as was demonstrated in Ref. Bulgac:2024a , even if the system is started in a fully canonical basis, in the subsequent time evolution the quasiparticle states are not anymore canonical states, and at each time step one has to identify the new canonical set of wave functions. Only in this case one can clearly see that the single odd quasiparticle state occupied with probability 1, ceases to maintain the same occupation probability. However, at each time $t>00$ there is only one canonical state occupied with exactly probability 1. We expect that this is a feature to be observed in any nuclear LACM of an odd fermion system, in either excited states of a nucleus or in heavy-ion collisions. A clear example when such a situation is encountered is when the angular momentum of the initial odd fermion is greater than $\approx A^{1/3}$, since such a state cannot exist with a significant occupation probability in either FF Bertsch:1997 ; Bulgac:2016 ; Bulgac:2019c ; Bulgac:2020 .
The funding for A.B. and M.K. from the Office of Science, Grant No. DE-FG02-97ER41014 and also the partial support provided by NNSA cooperative Agreement DE-NA0003841 is greatly appreciated. The work of I.A. and I.S. was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. The Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy Contract No. 89233218CNA000001. This research used resources of the Oak Ridge Leadership Computing Facility, which is a U.S. DOE Office of Science User Facility supported under Contract No. DE-AC05-00OR22725. The report is published as a preprint with numbers NT@UW-24-11 and LA-UR-24-xxxxxx.